Loc weight 14 pounds. The HCP prescribes an IV infusion of 5% dextrose in 1/4 normal saline to be infused at 8 ml / kg / HR. How many milliliters of the prescribed solution should the nurse Infuse each hour
Answer #1

Given:

Weight = 14 pounds = 6.363 (to convert ponds to kg divide by 2.2 )

Recommended prescribed dose = 8 ml / kg / hr

To find :

Dose of LOC= ?

Formula :

Dose to be given = weight * recommended prescribed dose

= 6.363 * 8 = 50.904 ml / hour

Therefore 50.904 ml of the prescribed solution should the nurse infuse each hour.